OpenStudy (anonymous):

verify 1-tanx ------- = cscx - secx sinx

OpenStudy (kainui):

Separate the fraction to be 1/sinx+tanx/sinx. Then separate out tanx into sinx/cosx Think about it for a little bit and try to understand why.

OpenStudy (anonymous):

{1/sin(x)}- { tan(x)/sin(x)}= csc(x)- { sin (x)/cos(x) *1/sin (x)}= csc(x)-i/cos(x)=csc(x)-sec(x)

OpenStudy (anonymous):

i still am confused

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

wait ok

OpenStudy (anonymous):